OpenStudy (anonymous):

3. Each of the folowing quadratic equation has two equal roots.Find the values of p. a)3x^2-px+12 pls help me to check whether my answer and working is correct or not. @ganeshie8

OpenStudy (anonymous):

\[x^2-\frac{ p }{ 3 }x+4=0\]

OpenStudy (anonymous):

therefore the roots are +-2

OpenStudy (anonymous):

\[(2)^2-\frac{ p }{ 3 }(2)+4=0\]

OpenStudy (anonymous):

or

OpenStudy (anonymous):

\[(-2)^2-\frac{ p }{ 3 }(-2)+4=0\]

OpenStudy (anonymous):

therefore, p=+-12

OpenStudy (anonymous):

is my answer correct?

OpenStudy (anonymous):

If you want the roots to be equal, make the discriminant 0.

OpenStudy (anonymous):

\[ (-p)^2-4(3)(12)=0 \]

OpenStudy (triciaal):

 do you know how to expand?